The phone system within a school assigns 4-digit phone numbers.
If the phone numbers are randomly assigned, what is the probability that you will be assigned a phone number that ends in 0, if the digits may repeat.

Answer by stanbon(75887)

The phone system within a school assigns 4-digit phone numbers.
If the phone numbers are randomly assigned, what is the probability that you will be assigned a phone number that ends in 0, if the digits may repeat.
-----
Ways to pick each of the first three digits: 10
Ways to pick the 4 digit: 1
----
# of phone numbers ending in zero: 10^3*1 = 1000
---
Total # of possible phone numbers: 10^4
---------------------
P(phone number ending in 0) = 1000/10,000 = 1/10
====================================================
Cheers,
Stan H